Sales of personal computers are expected to increase in the twenty first century. In 2000 and 2001, a combined total of 264 million personal computers were sold worldwide, with 12.9% more computers sold in 2001 than in 2000.How many personal computers were sold in 2000, and how many were sold in 2001? Round yours answers to the nearest million.
Answer by chana(69)

Sales of personal computers are expected to increase in the twenty first century. In 2000 and 2001, a combined total of 264 million personal computers were sold worldwide, with 12.9% more computers sold in 2001 than in 2000.How many personal computers were sold in 2000, and how many were sold in 2001? Round yours answers to the nearest million.
sol:
let x be the PC sold in 2000
y be the PC sold in 2001
x + y = 264 MILL
12.9% more computers sold in 2001 than in 2000
so y = (112.9/100)* x
x ( 1 + (112.9/100)*x) = 264 MILL
x = 124,001,879 PC sold